OpenStudy (anonymous):

Calculus 1: Show that the inequality is true.

OpenStudy (anonymous):

dude do your own work

OpenStudy (anonymous):

Show that: \[0 \le \int\limits_{2}^{3}\sin \sqrt[3]{x}dx \le 1\]